OpenStudy (andijo76):

Please HELP....If np>5 and nq>5 estimate P (more than 9) with n=12 and p=0.7 by using the normal distribution as an approximation to the binomial distribution if np <5 or nq<5 then state the normal approximation is not suitable

OpenStudy (anonymous):

hint: read the book :D

OpenStudy (andijo76):

q=1-p -1-0.7 -0.6 np-12*0.7=8.4 nq= 13*0.6= 7.2 np=5.2 nq=7.8 μ=np =12*0.7 =8.4 13∗0.4∗0.6−−−−−−−−−−√ =2.244 x=5-0.5 =4.5 x-u/o =(5-0.5)/-6=2.244 =0.6684 P(z<-0.6884)=

OpenStudy (andijo76):

now i am stuck

OpenStudy (andijo76):

i figured it out at least i hope i am right lol